[Brahms-l] Shift report 20040221 00:00-08:00

From: Apache <apache@bnl.gov>
Date: Sat Feb 21 2004 - 08:18:03 EST
------------------------------------------------------------
Date: Sat Feb 21 08:18:03 2004, Shift: 00:00-08:00
Supervisor: DO, with: 
Log book#15 pages 63 - 63

Shift Summary:
------------------------------------------------------------
1. One store from 1 to 5:30 am. New store started at 7am.

2. Data: runs 10457 -> 10460,  10464 still running
   FS  4d   B1/1: 576k tr2 
   MRS 40d  B1/2: 773k tr3

   TPC ped:     10462
   Non TPC ped: 10461

3. No problem until I turned TPC and DC's HV when beam
   conditions were ok (~7.30am). 
   TPM1 tripped, I got it back to
   business but then mainframe L1458B (DC's) was off.
   I think it is brahvo which screwed up the mainframe
   state. Then, all DC demand HV's were set to 0, could 
   not load config files due to miserable crashes, had 
   to set all demand HVs by hand for each DC channel! 
   and then switch them on by hand as well. 
   This of course caused a waste of time of ~ 1/2 hour 
   of good beam...
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sat Feb 21 08:18:18 2004

This archive was generated by hypermail 2.1.8 : Sat Feb 21 2004 - 08:18:34 EST